Lance Lopez went to his bank to find out how long it take for $2100 to amount to $2225 at 5% simple interest. Can you solve lances problem?
Answer #1

So, we have to you the time value of money fomula

FV = PV*(1+r)^n where FV = 2225, PV = 2100, r = 0.05 and n = ? where we have to calculate n

2225 = 2100*(1+0.05)^n

2225 =2100*1.05^n

2225/2100 = 1.05^n

1.05952381 = 1.05^n

Taking log both sides

log 1.05952381 = n* long 1.05

n = log 1.05952381/ log 1.05

n = 1.185 years

So, it takes 1.185 years for $2,100 to grow to $2,225